The equilibrium constant for the production of ammonia, used as fertilizer, is 3.65x 10

8 at 25 ⁰C. N2(g) + 3 H2(g) ⇋ 2 NH3(g) . Calculate the equilibrium concentration of hydrogen gas, if at

equilibrium the concentrations of nitrogen and ammonia are 2.56 M and 5.76 M


1
Expert's answer
2022-02-20T08:36:20-0500

"k_c=\frac{[NH_3]^2}{[N_2][H_2]^3}"


"[H_2]=\sqrt[3]{\frac{[NH_3]^2}{k_c[N_2]}}=\sqrt[3]{\frac{(5.76)^2}{3.65\times10^8\times2.56}}=0.00329\ M"


Answer: 0.00329 M